OpenStudy (anonymous):

Expand this expression, write your answer in the form of a + b\[\sqrt{c}\] where a, b, c are integers: 1) \[\sqrt{5}\] (2+\[\sqrt{5}\])?

OpenStudy (anonymous):

√5(2+√5)? - I mean this, sorry im new to the symbols on the site.

OpenStudy (anonymous):

It's not clear. You need to write it again.

OpenStudy (anonymous):

=\[2\sqrt{5}+25\]

OpenStudy (anonymous):

Hm, ah yes- thank you agdg :)- Is it possible to tell me how you got to that answer, I forgot the method on doing it, if its not too much hassle hehe :)

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

Simply multiple each term within the parentheses by \[\sqrt{5}\]

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Ah yes, but I dont understand why √5 X √5 = 25, what if it was √5 X √4 = 20?, Or am i wrong on that?

OpenStudy (anonymous):

\(\sqrt{5}(2+\sqrt{5})=2\sqrt{5}+5\).

OpenStudy (anonymous):

\(\sqrt{5}\times\sqrt{5}=5\), \(\sqrt{4}\times \sqrt{4}=4\)... and so on!